The amount of chocolates in a box from a particular production line is normally distributed with μ = 110 grams and σ = 25 grams. A sample of 25 boxes is to be selected.

What is the probability that the sample mean will be greater than 100 grams?
Answer #1

Solution:

Given data:

normally distributed with μ = 110 grams

σ = 25 grams

n =  25 boxes

Now we have to find out the :

The probability that the sample mean will be greater than 100 grams:

sample mean will be greater than 100 grams X > 100

probability that the sample mean will be greater than 100 = P( X > 100 )

Now, P( X > 100 ) = p( rac{x-mu }{sigma } > rac{x-mu}{sigma })

=   p( z > rac{x-mu}{sigma })

=  100- 110 2>

= ー10 p(z >·クー) 25

=   z-0.4)

We know, p( x>a ) = 1-p(x<a)

  z-0.4)  = I-p(z <-0.4)

From the z table we can get the value of - 0.4 is " 0.3446 "

= 1 - 0.3446

= 0.6554

herefore P( X > 100 ) =   0.6554

herefore Probability that the sample mean will be greater than 100 grams is " 0.6554 ".